# Automatically Interpreting Millions of Features in Large Language Models

### v2

@@ A.6.9 Dependence on SAE size and location
131k latents Gemma 2 9b 0.76 (0.67–0.86) 0.74 (0.63–0.85)
16k latents Gemma 2 9b 0.73 (0.63–0.83) 0.70 (0.59–0.79)
@@
262k latents Llama 3.1 8b MLP 0.79 (0.69–0.86) 0.79 (0.64–0.85)
262k latents Llama 3.1 8b 0.81 (0.71–0.86) 0.83 (0.68–0.85)
@@
A.7 Automatically interpreting interventions
A.7.1 Scoring implementation
For each latent, we sample a pool of 40 length-64 prompts from RPJv2 (Computer, 2023), of which 30 i.i.d. prompts are taken for scoring, while the remaining 10 are used by the explainer. The pool is sampled among nonzero activating contexts, stratified by the quintile of the latent’s max activation in that context (i.e., the nonzero activating documents are sorted by the context’s maximum activation of the current latent, then 8 contexts are sampled from the first quintile, 8 from the second etc.). Each of these prompts is then truncated to include only the first token on which the latent activates, so that all previous activations for the latent are 0. Activations on the first token position are ignored because the SAEs were not trained on that position.
We filter out latents that activates on less than 200 of the 10 million RPJ-v2 tokens.
We then sample generations with a maximum of 8 new tokens from the subject model. For generations with the intervention, we perform an additive intervention on the latent at all token positions after and including the final prompt token. We tune the intervention strength of each latent to various KL-divergence values on the scoring set with a binary search. We stop the binary search when the KL divergence is within 10% of the desired value [^3] Sometimes the KL-divergence is not perfectly monotonic in the intervention strength so 10% error is exceeded. We report the average KL divergence that we observe in Figure A6. For our zero-ablation experiment, instead of doing an additive intervention we clamp the latent’s activation to 0. Specifically, the hidden states are encoded by the SAE, then the SAE reconstruction error is computed using a clean decoding, then the SAE encoding is clamped and decoded, and the error is added to the clamped decoding.
The scorer is Llama-3.1 8B base. We use the base model for improved calibration, and prompt it as follows. {mdframed}
@@ A.7.2 Generating interpretations
We use the following prompt for the explainer, with 3 few-shot examples truncated for brevity. {mdframed}
@@
A.7.3 Baselines
Random SAE. We experiment with a random top-k SAE with $k=50$, roughly the average sparsity of the gemma SAEs. The encoder is initialized with 131,072 spherically uniform unit-norm latents, and the decoder is initialized to its transpose. We use a random SAE with TopK activations because we need sparsity for our sampling procedure to work properly.
Random interpretations. For each layer and target KL value, we compute a random interpretation baseline where we shuffle the interpretations across latents.
A.7.4 Intervention interpretability results
@@
Figure A6: Average intervention score vs layer. SAE latents in later layers have more explainable effects on output. Random latents, however, have uninterpretable effects on output, even at late layers.

### v3

@@ F.8 Dependence on SAE size and location
131k features Gemma 2 9b 0.76 (0.67–0.86) 0.74 (0.63–0.85)
16k features Gemma 2 9b 0.73 (0.63–0.83) 0.70 (0.59–0.79)
@@
262k features Llama 3.1 8b MLP 0.79 (0.69–0.86) 0.79 (0.64–0.85)
262k features Llama 3.1 8b 0.81 (0.71–0.86) 0.83 (0.68–0.85)
@@
Appendix G Automatically interpreting interventions
G.1 Scoring implementation
For each feature, we sample a pool of 40 length-64 prompts from RPJv2 (Computer, 2023), of which 30 i.i.d. prompts are taken for scoring, while the remaining 10 are used by the explainer. The pool is sampled among nonzero activating contexts, stratified by the quintile of the feature’s max activation in that context (i.e., the nonzero activating documents are sorted by the context’s maximum activation of the current feature, then 8 contexts are sampled from the first quintile, 8 from the second etc.). Each of these prompts is then truncated to include only the first token on which the feature activates, so that all previous activations for the feature are 0. Activations on the first token position are ignored because the SAEs were not trained on that position.
We filter out features that activates on less than 200 of the 10 million RPJ-v2 tokens.
We then sample generations with a maximum of 8 new tokens from the subject model. For generations with the intervention, we perform an additive intervention on the feature at all token positions after and including the final prompt token. We tune the intervention strength of each feature to various KL-divergence values on the scoring set with a binary search. We stop the binary search when the KL divergence is within 10% of the desired value [^4] Sometimes the KL-divergence is not perfectly monotonic in the intervention strength so 10% error is exceeded. We report the average KL divergence that we observe in Figure A6. For our zero-ablation experiment, instead of doing an additive intervention we clamp the feature’s activation to 0. Specifically, the hidden states are encoded by the SAE, then the SAE reconstruction error is computed using a clean decoding, then the SAE encoding is clamped and decoded, and the error is added to the clamped decoding.
The scorer is Llama-3.1 8B base. We use the base model for improved calibration, and prompt it as follows.
@@ G.2 Generating interpretations
We use the following prompt for the explainer, with 3 few-shot examples truncated for brevity.
@@
G.3 Baselines
Random SAE. We experiment with a random top-k SAE with $k=50$, roughly the average sparsity of the gemma SAEs. The encoder is initialized with 131,072 spherically uniform unit-norm features, and the decoder is initialized to its transpose. We use a random SAE with TopK activations because we need sparsity for our sampling procedure to work properly.
Random interpretations. For each layer and target KL value, we compute a random interpretation baseline where we shuffle the interpretations across features.
G.4 Intervention interpretability results
@@
Figure A6: Average intervention score vs layer. SAE features in later layers have more explainable effects on output. Random features, however, have uninterpretable effects on output, even at late layers.